OpenStudy (anonymous):

Fill in the missing number to complete the factorization: x^3+14x^2+59x+70=(x+2)(x+5)(x+__________)

OpenStudy (anonymous):

@amistre64 @ganeshie8 @Directrix

OpenStudy (anonymous):

@zepdrix

ganeshie8 (ganeshie8):

Hint : constant term

zepdrix (zepdrix):

\[\large\rm x^3+14x^2+59x+\color{orangered}{70}=(x+\color{orangered}{2})(x+\color{orangered}{5})(x+\color{orangered}{?})\]

OpenStudy (jhannybean):

Guessing the rational root theorem?

OpenStudy (anonymous):

I don't get what you mean by constant term

OpenStudy (anonymous):

The Remainder Theorem

OpenStudy (anonymous):

with Factorization

ganeshie8 (ganeshie8):

constant term is the one without any x's attached to it

OpenStudy (anonymous):

10?

ganeshie8 (ganeshie8):

\[\large\rm x^3+14x^2+59x+\color{orangered}{70}=(x+\color{orangered}{2})(x+\color{orangered}{5})(x+\color{orangered}{?})\] look at left hand side, whats the constant term ?

OpenStudy (anonymous):

7?

ganeshie8 (ganeshie8):

do you mean 70 ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so 70 would be the constant term for the (x+____)

ganeshie8 (ganeshie8):

since above is an equation, left side and right side must have the same constant term

ganeshie8 (ganeshie8):

\[\large\rm x^3+14x^2+59x+\color{orangered}{70}=(x+\color{orangered}{2})(x+\color{orangered}{5})(x+\color{orangered}{?})\] look at right hand side, what could be the constant term ?

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

you get the constant term on right hand side by multiplying these : \(\color{red}{2,5,?}\)

ganeshie8 (ganeshie8):

\[\color{Red}{70 = 2*5*?} \implies \color{red}{\dfrac{70}{10} = ?}\]

OpenStudy (anonymous):

I DID SAID 7

ganeshie8 (ganeshie8):

you're correct